OpenStudy (anonymous):

PLEASE HELP :( What is the equation of the line whose y-intercept is 3 and slope is 1? A. y = x - 3 B. y = x + 3 C. y = 3x + 1

OpenStudy (anonymous):

b

OpenStudy (anonymous):

y = mx + c Where m is the gradient and c is the y intercept . Therefore y = x + 3

OpenStudy (anonymous):

y = 2x + 6 has slope 2, therefore any line perpendicular to it will have slope -1/2 (the inverse reciprocal). The point (0,3) lies on the y-axis and is therefore the desired line's y-intercept. Since you need the equation in slope-intercept form, you've just got to throw the numbers into the equation. y = mx + b m = -1/2 b = 3 Therefore, DRUM ROLL PLEASE, y = (-1/2)x + 3 Hope this helped!